OpenStudy (anonymous):

Find the volume of a can of soup that has 16cm height and a radius of 5 cm use 3.14 as pi

OpenStudy (anonymous):

V=πr2h is the equation try plugging in your height and radius

OpenStudy (dayakar):



OpenStudy (jmark):

Can is of cylindrical shape. Volume = pi.r^2.h = 3.14*16*5*5 = 1256 cubic cm Follow TutorVista at https://twitter.com/tutorvista